how many grams of oxygen are needed to completely burn 35.6 grams of C3H8?
1
Expert's answer
2012-12-03T10:26:45-0500
The molar mass of C3H8 is
M(C3H8) = 3*12g/mol + 8*1g/mol = 36 + 8 = 44 g/mol Theamount of C3H8 is
n(C3H8) = m(C3H8) / M(C3H8) = 35.6 g / 44 g/mol = 0.81mol The reaction of burning of C3H8 is
C3H8 + 5O2 = 3CO2 + 4H2O
According to reaction the amount of oxygen is
n(O2) = 5*n(C3H8) = 5* 0.81 = 4.05 mol
The mass of oxygen is
m(O2) = n(O2) * M(O2) = 4.05 mol * 32 g/mol = 129.6 g